OpenStudy (gabylovesyou):

@hartnn Two Olympic swimmers, Ryan and Michael, create unique ripple waves in the water. The function created by Ryan is f(x) = 4x2 + 7. The function created by Michael is f(x) = 4x2 - 5x + 3. Describe to Ryan and Michael, using complete sentences, which of the operations—addition, subtraction, multiplication, and division—will result in the largest degree function and which operation will result in the smallest degree function.

hartnn (hartnn):

do you know what 'degree of a function' means ?

OpenStudy (gabylovesyou):

yes sir :)

hartnn (hartnn):

then do you know what operation can increase the degree?

OpenStudy (gabylovesyou):

use an operation for both functions.. for example add f(x) = 4x^2 + 7 and f(x) = 4x^2- 5x + 3 ?

OpenStudy (gabylovesyou):

thats what it means ?

hartnn (hartnn):

what is the degree of 'x^2' ? what is the degree of 'x^3' ? what is x^2 * x^3 ?? what is the degree of the above result ?

OpenStudy (gabylovesyou):

6

hartnn (hartnn):

sure? x^2 * x^3 = x^{2+3} = .. ?

OpenStudy (gabylovesyou):

oh right cuz u add em cx so 5

hartnn (hartnn):

yes, so can we conclude multiplying polynomial increases the degree? :)

OpenStudy (gabylovesyou):

yes

hartnn (hartnn):

so if multiplication increases, what operation will decrease the degree?

OpenStudy (gabylovesyou):

subtraction o-o ?

hartnn (hartnn):

no...what is the opposite operation of multiplication ?

OpenStudy (gabylovesyou):

division

hartnn (hartnn):

yes. though in some cases (like this), subtraction will also result in reduction of degree :)

OpenStudy (gabylovesyou):

how would i do it with these functions ?

OpenStudy (gabylovesyou):

what i mean is use multiplication and division with these functions given

hartnn (hartnn):

you need to do all those 4 operations. check the result. and based on that, prove that multiplication will increase the degree subtraction and division will reduce the degree and addition will have no effect on degree

OpenStudy (gabylovesyou):

but how do i do that.. 4x^2 * 4x^2 * x^1 ? or how

hartnn (hartnn):

you can only describe by giving this example \(Multiplication : x^2 \times x^2 = x^4 \) degree is maximum \(Subtraction : x^2 -x^2 = 0 \) degree is minimum

OpenStudy (gabylovesyou):

ok ...

OpenStudy (gabylovesyou):

x^2 + x^2 = x^4 ?

hartnn (hartnn):

x^2+ x^2 = 2x^2 same degree

OpenStudy (gabylovesyou):

x^2 / x^2 = x ?

hartnn (hartnn):

x^2/x^2 = 1 degree = 0

OpenStudy (gabylovesyou):

wait but subtraction and division both r = to 0 degree ?

hartnn (hartnn):

yes, division always reduces the degree but subtraction reduces sometimes only... when the co-efficient of the highest degree term is same

OpenStudy (gabylovesyou):

in this case.. it reduced in both.. so which is least ?

hartnn (hartnn):

in this particular example, both gives you the least :) you can go with subtraction, because it is easy to calculate :)